OpenStudy (anonymous):

Simplify using the order of operations: 6(-9) – (-8) – (-7)(4) + 11

OpenStudy (anonymous):

first you need to know the order of operations: PEMDAS. Paranthesis, Exponents, Multiplication, Division, Addition, Subtraction.

OpenStudy (anonymous):

-3 + 8 - -28 + 11

OpenStudy (anonymous):

-3 + 8 + 28+11 = 45

OpenStudy (anonymous):

check it with a calculator i added/subtracted in my head. Remember multiplication before addition/subtraction.

OpenStudy (anonymous):

PEMDAS 6(-9) – (-8) – (-7)(4) + 11 =-54+8-(-28)+11 =-54+8+28+11 =-7

OpenStudy (anonymous):

i realized theres multiplication at the beginning so I am wrong. let me redo that.

OpenStudy (gw2011):

-54+8-(-28)+11 -54+8+28+11=-7

OpenStudy (anonymous):

Answer is -7 with a calculator. I forgot the first multiplication part.

OpenStudy (anonymous):

Parentheses/Grouping Exponents Multiplication Division (from left to right) Addition Subtraction (from left to right)